# TReMu: Towards Neuro-Symbolic Temporal Reasoning for LLM-Agents with Memory in Multi-Session Dialogues

### v1

@@ 4.3 Execution Failure Study
We observe that LLMs occasionally generate Python code with syntax errors, such as typos or mismatched brackets. To delve into this, we measure the percentage of generated code that fails to execute and, during inference, we regenerate the code when such errors occur. The results, shown in Figure 2, indicate that the percentages of execution failure are generally low across all three LLMs, demonstrating the reliability of our Python-based symbolic reasoning approach. As expected, GPT-4o exhibits the lowest rate of execution failure, while GPT-3.5-Turbo has the highest, corresponding to the overall performance differences we demonstrate above in temporal reasoning among these models. This likely reflects the inherent performance gap between the LLMs.
@@ 4.4 Case Study
In Figure 3, with CoT, even though GPT-4o successfully identifies that the key temporal information is "Sharon’s survival course started on 12 March 2020," but it gets confused with "week-long course" and infers the end date of the course, incorrectly selecting "Unanswerable." In contrast, with our framework’s time-aware memorization, the model retrieves the event from memory along with its properly inferred time. During the reasoning stage, the model utilizes this memory to distinguish between when the speaker, Sharon, mentioned the event (03/16/2020) and when the event occurred (03/12/2020). Then the model defines the corresponding variable in the generated code, i.e., t__\__start__\__course, to precisely capture the time.
@@ 5 Related Work
Several studies have focused on multi-session dialogues and developed benchmarks to address this task. Xu et al. introduced MSC, the first multi-session dataset incorporating time intervals between sessions. Similarly, Bae et al. proposed a dynamic memory management method to maintain up-to-date user information and introduced a Korean multi-session dialogue dataset. Jang et al. created the CONVERSATION CHRONICLES dataset, designed for long-term conversations that integrate time intervals and detailed speaker relationships. More recently, Maharana et al. introduced LoCoMo, a dataset featuring long-term dialogues spanning up to 35 sessions. While our work is situated within this context, it specifically targets temporal reasoning, addressing the temporal characteristics of relative time and cross-session dependency, which have not been explicitly explored in prior research.
@@ 6 Conclusion
In this paper, we address the critical challenge of temporal reasoning in multi-session dialogues for LLM-agents. We present a method to construct a temporal reasoning evaluation benchmark by augmenting dialogues from LoCoMo, specifically targeting the temporal characteristics of relative time and cross-session dependency. Furthermore, we introduce a novel framework which combines time-aware memorization through timeline summarization with neuro-symbolic temporal reasoning by translating temporal questions into executable Python code. Through extensive evaluations, we demonstrate that our benchmark presents significant challenges, and our framework substantially outperforms baseline methods in improving temporal reasoning for multi-session dialogues.
@@ 7 Limitations
Our work has several limitations. First, our constructed temporal reasoning benchmark consists of multiple-choice questions, while real multi-session dialogues involve users interacting with LLM agents in a generative manner. A benchmark that incorporates generative temporal reasoning questions could more accurately reflect the agents’ reasoning abilities in practical scenarios. We plan to extend our current benchmark to support generative question-answering in future work.
Additionally, we observe that the LoCoMo dialogues contain some temporal ambiguities that can affect reasoning over related events. For instance, if a speaker mentions on a Tuesday that an event occurred "last Monday," it is unclear whether the event happened the previous day or on the Monday of the prior week. However, in real applications, LLM agents could potentially resolve such ambiguities by asking follow-up clarifying questions.

### v2

@@ 4.4 Execution Failure Study
We also measure the percentage of generated code that fails to execute and, during inference, we regenerate the code when such errors occur. The results, shown in Figure 2, indicate that the percentages of execution failure are generally low across all three LLMs, demonstrating the reliability of our Python-based symbolic reasoning approach. As expected, GPT-4o exhibits the lowest rate of execution failure, while GPT-3.5-Turbo has the highest, corresponding to the overall performance differences we demonstrate above in temporal reasoning among these models. This likely reflects the inherent performance gap between the LLMs.
@@ 4.5 Case Study
In Figure 3, with CoT, even though GPT-4o successfully identifies that the key temporal information is "Sharon’s survival course started on 12 March 2020," but it gets confused with "week-long course" and infers the end date of the course, incorrectly selecting "Unanswerable." In contrast, with our framework’s time-aware memorization, the model retrieves the event from memory along with its properly inferred time. During the reasoning stage, the model utilizes this memory to distinguish between when the speaker, Sharon, mentioned the event (03/16/2020) and when the event occurred (03/12/2020). Then the model defines the corresponding variable in the generated code, i.e., t_\_start_\_course, to precisely capture the time.
@@ 5 Related Work
Except from the massive research efforts on common dialogue generation Ge et al. (2023b), several studies have developed multi-session dialogues benchmarks. Xu et al. introduced MSC, the first multi-session dataset incorporating time intervals between sessions. Similarly, Bae et al. proposed a dynamic memory management method to maintain up-to-date user information and introduced a Korean multi-session dialogue dataset. Jang et al. created the CONVERSATION CHRONICLES dataset, designed for long-term conversations that integrate time intervals and detailed speaker relationships. More recently, Maharana et al. introduced LoCoMo, a dataset featuring long-term and multi-modal dialogues. While our work is situated within this context, it specifically targets temporal reasoning, addressing the temporal characteristics of relative time and cross-session dependency, which have not been explicitly explored in prior research.
@@ 6 Conclusion
In this paper, we address the critical challenge of temporal reasoning in multi-session dialogues for LLM-agents. We present a method to construct a temporal reasoning evaluation benchmark by augmenting dialogues from LoCoMo and covering the temporal characteristics of relative time and cross-session dependency. Furthermore, we introduce a novel framework which combines time-aware memorization through timeline summarization with neuro-symbolic temporal reasoning by translating temporal questions into executable Python code. Through extensive evaluations, we demonstrate that our benchmark presents significant challenges, and our framework substantially outperforms baseline methods in improving temporal reasoning for multi-session dialogues.
@@ 7 Limitations
Our work has several limitations:
Assessment in QA settings. Our evaluation follows the standard practice of recent temporal reasoning benchmarks like Zhou et al. (2019); Wang and Zhao (2024), using a multiple-choice format for reliable assessment. A more ideal setting would be evaluating in generative dialogue settings. However, generative QA evaluations pose significant challenges due to variations in temporal expressions (e.g., "January 1st" vs. "01/01"), making exact match and F1 token score unreliable. Thus, we prioritize benchmark reliability and accuracy over a more ambitious generative QA setting. We leave the extension of the current benchmark to generative dialogue settings as future work
Open-sourced Models. Our current experiments mainly focus on close-sourced models. However, as we have pointed out in our Experiment section $\S$ 4.1, we found most open-source LLMs cannot handle the long dialogue inputs from LoCoMo, for example only about 10% of dialogues can be fed into Llama-3-70B. And even for those shorter dialogues that can be fed into Llama-3-70B, we notice that the model gets lost and fails to follow instructions, even failing to generate in the desired format. We therefore consider the adaptation of our framework to open-sourced models as future work. Note that we will release our code and dataset for reproducibility.